z=e^(x-2y) x=cost y=sint dz/dt=?


1
Expert's answer
2021-04-13T16:28:14-0400

Given: "z=e^{x-2y},x=cost,y=sint"

Require to find "\frac{dz}{dt}"


Recollect the following Chain Rule:


If "z=f(x,y),x=g(t),y=h(t)" , then "\frac{dz}{dt}=\frac{\partial z}{\partial x}\frac{dx}{dt}+\frac{\partial z}{\partial y}\frac{dy}{dt}"


Now "z=e^{x-2y}\Rightarrow \frac{\partial z}{\partial x}=e^{x-2y}(1-2(0))=e^{x-2y}" , "\frac{\partial z}{\partial y}=e^{x-2y}(0-2(1))=-2e^{x-2y}"


And "x=cost,y=sint\Rightarrow \frac{dx}{dt}=-sint, \frac{dy}{dt}=cost"


Using the Chain Rule, we get

"\frac{dz}{dt}=\frac{\partial z}{\partial x}\frac{dx}{dt}+\frac{\partial z}{\partial y}\frac{dy}{dt}"

"\Rightarrow \frac{dz}{dt}=(e^{x-2y})(-sint)+(-2e^{x-2y})(cost)"

"\Rightarrow \frac{dz}{dt}=-sinte^{x-2y}-2coste^{x-2y}"

"\Rightarrow \frac{dz}{dt}=-e^{x-2y}[sint+2cost]"


Therefore "\Rightarrow \frac{dz}{dt}=-(sint+2cost)e^{x-2y}"